# f(R) gravity with spacetime torsion

### v1

@@ II Ricci Tensor with torsion
where the lower and upper indices are with respect to $\widetilde{g}_{ab}$. The above action resembles with a scalar-tensor action along with a 2-rank massless antisymmetric tensor field ($Z_{[ab]}$), where the scalar field gets non-minimally coupled with the 2-rank tensor field. Note that $\phi(x)$ acts as a scalar field with the potential $V(A(\phi))=\frac{Af^{\prime}(A)-f(A)}{f^{\prime}(A)^{2}}$, and the antisymmtric tensor field carries the signature of the spacetime torsion. Therefore in the context of $f(R)$ gravity in presence of spacetime torsion, the higher curvature d.o.f manifests itself as a scalar field d.o.f which couples with the massless antisymmetric tensor field (or equivalently, with the torsion field) through derivative kind of coupling. It is important to note that the scalar field and the torsion do not propagate independently, actually the torsion field gets a source term from $\phi(x)$. Clearly such coupling between the scalar and the torsion fields introduces new three or four point vertices which may have interesting phenomenological implications both in cosmology as well as in particle physics. Here it is important to note that the 3-point interaction vertex between the $\phi(x)$ and the $Z_{[ab]}$ contain the factor $\kappa$, while the 4-point vertex gets suppressed by $\kappa^{2}$ (this can also be understood from dimensional analysis). Therefore the interaction that can give the most significant effects is given by the terms in the second line of action (III), i.e., the 3-point interaction between $\phi$ and $Z_{[ab]}$. Being a derivative coupling, it should have significant effects during the early universe cosmological phenomena where the energy scale of the universe is of $\sim 10^{-3}$ order of the Planck scale. Here we should mention that an equivalent scalar-tensor representation of Cartan f(R) theory has been demonstrated in Inagaki:2022blm where the authors showed that, similar to normal f(R) gravity, Cartan f(R) theory can also be represented to a minimally coupled scalar-tensor theory (without any antisymmetric tensor field). Actually the formalism of the current work is different than Inagaki:2022blm , in particular, we transform the “f(R) gravity with torsion” from Jordan to Einstein frame based on a conformal transformation of metric, unlike to Inagaki:2022blm where the authors rewrite the Cartan f(R) theory in Jordan frame itself by a different fashion without considering any conformal transformation of the metric. This makes our present scenario essentially different from the earlier ones.
The massless rank-2 antisymmetric tensor field may be identified with the Kalb-Ramond (KR) field. The minimally coupled scalar-tensor theory along with a KR field (i.e., where there is no coupling between the scalar field and the KR field) proves to be useful for explaining various cosmic phenomena during early universe, as showed by some of our authors in Elizalde:2018rmz ; Paul:2022mup . In particular, the Kalb-Ramond is able to trigger a bouncing universe or an inflationary universe depending on the initial conditions (see Elizalde:2018rmz ; Paul:2022mup ). Regarding the inflation, it turns out that the KR field energy density decays at a faster rate compared to the radiation and the pressureless dust with the expansion of the universe — which indicates that it has significant effects during the early universe, in particular, it enhances the tensor perturbation amplitude and hence the tensor-to-scalar ratio. Consequently the KR field gives well contributions on enhancing primordial gravitational waves generated from Bunch-Davies vacuum during the early universe. Although the KR field has significant effects during the early universe, at the same time, it is also important to realize that the $present$ stage of the universe carries practically no observable footprints of higher rank antisymmetric tensor fields including the Kalb-Ramond one. Thus a natural question arises from observable signatures of scalar, fermion, and vector d.o.f. in our Universe along with spin 2 symmetric tensor field in the form of gravity: why is our Universe is free of any perceptible signature of massless antisymmetric tensor modes? A possible explanation of this issue can be found from the minimally coupled scalar-tensor-KR theory Elizalde:2018rmz ; Elizalde:2018now ; Paul:2020duu ; Paul:2022mup .
Thus as a whole, the Einstein frame action of $f(R)$ gravity with spacetime torsion is given by a non-minimally coupled scalar-tensor theory with a rank-2 massless antisymmetric tensor field, in particular, by the action (III) . The intriguing effects of the minimally coupled scalar-tensor-KR theory, as shown in Elizalde:2018rmz ; Elizalde:2018now ; Paul:2020duu ; Paul:2022mup , points the larger importance of the action (III) where the scalar field gets coupled with the torsion field. Such possibilities are expected to study, in detail, in some future work.
IV Summary
@@
In summary, we address the correspondence between the Jordan and the Einstein frame $f(R)$ gravity in presence of spacetime torsion. Without any torsion, it is well known that $f(R)$ gravity action can be mapped to a $minimally$ coupled scalar-tensor theory in the Einstein frame, where the scalar field appears with a potential that depends on the form of $f(R)$ under consideration. However the scenario changes with spacetime torsion, in particular, if the affine connection is considered to have an antisymmetric part. It turns out that after the inclusion of torsion, the $f(R)$ action can be recast to a $non-minimally$ coupled scalar-tensor theory with a 2-rank massless antisymmetric tensor field in the Einstein frame, where the scalar field gets coupled with the antisymmetric field. Actually the antisymmetric field carries the signature of the spacetime torsion. Therefore, interestingly, the scalar field (coming from the higher curvature d.o.f) and the torsion field do not propagate independently, actually the torsion field gets a source term from the scalar field. Furthermore, such interaction between the scalar and the torsion fields comes with a derivative coupling, and thus introduces a momentum dependent interaction vertex factor — which may have interesting phenomenological implications both in cosmology as well as in particle physics. In this regard, it may be mentioned that the 3-point interaction vertex between the $\phi(x)$ and the $Z_{[ab]}$ contain the factor $\kappa$, while the 4-point vertex gets suppressed by $\kappa^{2}$ (this can also be understood from dimensional analysis). Therefore the interaction that can give the most significant effects is given by the 3-point interaction between $\phi$ and $Z_{[ab]}$..
The rank-2 massless antisymmetric tensor field may be identified with Kalb-Ramond field which was studied a lot in the sector of cosmology, black holes and particle physics. Such intriguing effects of the KR field points the larger importance of $f(R)$ theory with spacetime torsion, where the torsion field shows a non-minimal coupling with the scalaron field. This is expected to study in some future work.

### v2

@@ 3 $f(R)$ action with torsion in Einstein frame
where the lower and upper indices are with respect to $\widetilde{g}_{ab}$, and $\widetilde{\bar{R}}$ is the Ricci scalar in Einstein frame formed by the symmetric affine connection. The above action resembles with a scalar-tensor action along with a 2-rank massless antisymmetric tensor field ($Z_{[ab]}$), where the scalar field gets non-minimally coupled with the 2-rank tensor field. Note that $\phi(x)$ acts as a scalar field with the potential $V(A(\phi))=\frac{Af^{\prime}(A)-f(A)}{f^{\prime}(A)^{2}}$, and the antisymmetric tensor field carries the signature of the spacetime torsion. Therefore in the context of $f(R)$ gravity in presence of spacetime torsion, the higher curvature d.o.f manifests itself as a scalar field d.o.f which couples with the massless antisymmetric tensor field (or equivalently, with the torsion field) through derivative kind of coupling. It is important to note that the scalar field and the torsion do not propagate independently, actually the torsion field gets a source term from $\phi(x)$. Clearly such coupling between the scalar and the torsion fields introduces new three or four point vertices which may have interesting phenomenological implications both in cosmology as well as in particle physics. Here it is important to note that the 3-point interaction vertex between the $\phi(x)$ and the $Z_{[ab]}$ contain the factor $\kappa$, while the 4-point vertex gets suppressed by $\kappa^{2}$ (this can also be understood from dimensional analysis). Therefore the interaction that can give the most significant effects is given by the 3-point interaction between $\phi$ and $Z_{[ab]}$. Being a derivative coupling, it is supposed to have significant effects during the early universe cosmological phenomena where the energy scale of the universe is of $\sim 10^{-3}$ order of the Planck scale. In order to demonstrate the cosmological implications of the interaction terms, let us go through the Friedmann equations for the Einstein action (27). For this purpose, we consider a spatially flat FLRW metric in the Einstein frame, as follows:
$d\tilde{s}^{2}=-dt^{2}+a^{2}(t)d\textbf{x}^{2}~{}~{},$ (28)
where $t$ is the cosmic time and $a(t)$ represents the scale factor of the universe. With the above metric along with the consideration that $\phi$ and $Z_{ab}$ depend only on the cosmic time, the corresponding Friedmann equations turn out to be (see the Appendix for the detailed derivation),
$\frac{3}{2\kappa^{2}}\bigg{(}\frac{\dot{a}}{a}\bigg{)}^{2}$ $=$ $\frac{1}{4}(\partial_{t}\phi)^{2}+\frac{V(\phi)}{2}$ (29)
-- $\frac{1}{3}\partial_{t}Z_{ad}\partial^{t}Z^{ad}-\frac{5\kappa}{6}\partial_{t}\phi Z_{ad}\partial^{t}Z^{ad}$
$+$ $\frac{5\kappa^{2}}{12}\partial_{t}\phi Z_{ad}\partial^{t}\phi Z^{ad}$
and
$\frac{3}{\kappa^{2}}\left[\frac{\ddot{a}}{a}+\bigg{(}\frac{\dot{a}}{a}\bigg{)}^{2}\right]$ $=$ $2V(\phi)-\frac{1}{2}(\partial_{t}\phi)^{2}$ (30)
-- $\frac{1}{2}\partial_{t}Z_{ij}\partial^{t}Z^{ij}+\kappa\partial_{t}\phi Z_{ij}\partial^{t}Z^{ij}$
-- $\frac{\kappa^{2}}{2}\partial_{t}\phi Z_{ij}\partial^{t}\phi Z^{ij}$
respectively, where an overdot symbolizes $\frac{d}{dt}$. Clearly the usual cosmological field equations for a scalar-tensor theory can be recovered from the above two in absence of the antisymmetric tensor field. It is evident that the antisymmetric tensor field couples with the scalar field through its velocity term, in particular, with $\dot{\phi}$. Such a coupling should have significant effects during the early stage of the universe (particularly before the BBN), here we list some of the possibilities:
  • In a typical inflationary context, at the beginning when the scalar field follows the slow roll condition(s), $\dot{\phi}$ remains negligible and hence the interaction of the antisymmetric field with the scalar field is not that significant. In that case, the Hubble parameter is dominated solely by the potential term $V(\phi)$ that leads to a de-Sitter inflation. However near the end of inflation the scalar field gets its acceleration, and as a result, the coupling between the antisymmetric and the scalar fields becomes considerable through $\dot{\phi}$, which in turn affects the end stage of inflation. In particular, Eq. (29) and Eq. (30) indicate that the antisymmetric field produces a deceleration effect on the cosmic expansion of the universe (this can be realized as Eq. (29) and Eq. (30) lead to $\ddot{a}\propto-(\dot{Z}_{ab})^{2}$ in absence of the scalar field), and thus the presence of $Z_{ab}$ along with its couplings with $\phi$ makes the inflation longer compared to the case where the antisymmetric field is absent.
  • After the end of inflation, the universe enters to a reheating stage when the scalar field generally decays to other forms of matter fields through its couplings. Therefore due to presence of the interaction terms between $\phi$ and $Z_{ab}$, the reheating scenario in the present context also results to the production of $Z_{ab}$ along with the radiation fluid from the scalar field energy density. Thus the universe, from the end of the reheating to the BBN $\sim 10^{-2}\mathrm{GeV}$, passes through an era dominated by the antisymmetric tensor field and the radiation fluid. Such a cosmological era should have interesting effects on the propagation of primordial GWs and even to the dark matter scenario where the antisymmetric field may act as the dark matter candidate.
  • The massless rank-2 antisymmetric tensor field may be identified with the Kalb-Ramond (KR) field. The minimally coupled scalar-tensor theory along with a KR field (i.e., where there is no coupling between the scalar field and the KR field) proves to be useful for explaining various cosmic phenomena during early universe, as showed by some of our authors in [12, 15]. In particular, the Kalb-Ramond is able to trigger a bouncing universe or an inflationary universe depending on the initial conditions (see [12, 15]). The intriguing effects of the minimally coupled scalar-tensor-KR theory, as shown in [12, 13, 14, 15], points the larger importance of the action (27) where the scalar field gets coupled with the torsion field.
These indicate the immense interest of the action in Eq. (27) in the cosmological sector starting from inflation to reheating stage. However the detailed study of the possibilities mentioned above are outside the focus of the present paper, and are expected to study in some near future works.
Thus as a whole, “$f(R)$ gravity with a non-zero antisymmetric part of affine connection” in the Jordan frame is dual to an action in Einstein frame represented by a “non-minimally coupled scalar-tensor + a massless rank-2 antisymmetric tensor field theory with totally symmetric affine connection”. It is well known that a normal $f(R)$ theory, in absence of any antisymmetric part of the affine connection, is dual to a minimally coupled scalar-tensor theory in the Einstein frame where the scalar field potential is fixed by the form of $f(R)$ under consideration. However in the present work, we find that the scenario gets different when the affine connection in the Jordan frame has an antisymmetric part which may be identified with the spacetime torsion, i.e. in the context of an Einstein-Cartan $f(R)$ theory. It turns out that $f(R)$ gravity with spacetime torsion in Jordan frame can be recast to a non-minimally coupled scalar-tensor theory with a rank-2 massless antisymmetric tensor field and with symmetric affine connection in the Einstein frame, in particular, by the action (27) where the antisymmetric tensor field couples with the scalar field through derivative coupling. Therefore we may argue that the extra degrees of freedom arising from the antisymmetric part of the affine connection in the Jordan frame shows as a massless rank-2 antisymmetric tensor field in the Einstein frame. This is a new result that generalizes the duality between Jordan and Einstein frames in presence of spacetime torsion.
4 Summary
@@
In summary, we address the correspondence between the Jordan and the Einstein frame $f(R)$ gravity in presence of spacetime torsion. Without any torsion, it is well known that $f(R)$ gravity action can be mapped to a $minimally$ coupled scalar-tensor theory in the Einstein frame, where the scalar field appears with a potential that depends on the form of $f(R)$ under consideration. However the scenario changes with spacetime torsion, in particular, if the affine connection is considered to have an antisymmetric part. It turns out that after the inclusion of torsion, the $f(R)$ action can be recast to a $non-minimally$ coupled scalar-tensor theory with a 2-rank massless antisymmetric tensor field in the Einstein frame, where the scalar field gets coupled with the antisymmetric field. Actually the antisymmetric field carries the signature of the spacetime torsion. Therefore, interestingly, the scalar field (coming from the higher curvature d.o.f) and the torsion field do not propagate independently, actually the torsion field gets a source term from the scalar field. Furthermore, such interaction between the scalar and the torsion fields comes with a derivative coupling, and thus introduces a momentum dependent interaction vertex factor — which may have interesting phenomenological implications both in cosmology as well as in particle physics. In this regard, it may be mentioned that the 3-point interaction vertex between the $\phi(x)$ and the $Z_{[ab]}$ contain the factor $\kappa$, while the 4-point vertex gets suppressed by $\kappa^{2}$ (this can also be understood from dimensional analysis). Therefore the interaction that can give the most significant effects is given by the 3-point interaction between $\phi$ and $Z_{[ab]}$.
Thus as a whole, the Einstein frame action of $f(R)$ gravity with spacetime torsion is given by a non-minimally coupled scalar-tensor theory with a rank-2 massless antisymmetric tensor field and with a totally symmetric affine connection, in particular, by the action (27) where the antisymmetric tensor field and the scalar field couple through derivative coupling. Such interactions should have some interesting cosmological implications (some of the possibilities are discussed after Eq. (30)) that are expected to study in some future work.
5 Appendix
As the variation of the action (27) with respect to the metric $\tilde{g}^{\mu\nu}$ (let us do it term by term):
Variation of the first term of the action (27):
$\delta(\sqrt{-\widetilde{g}}\widetilde{R}_{\mu\nu}\widetilde{g}^{\mu\nu})=\left(\widetilde{R}_{\mu\nu}-\frac{1}{2}\widetilde{g}_{\mu\nu}\widetilde{R}\right)\delta\widetilde{g}^{\mu\nu}\sqrt{-\widetilde{g}}~{}~{}.$
Variation of the second and third terms of the action (27):
$\delta\Big{(}\sqrt{-\widetilde{g}}\frac{1}{2}\widetilde{g}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-V(\phi)\Big{)}=\sqrt{-\widetilde{g}}\Big{[}\frac{1}{2}\partial_{\mu}\phi\partial_{\nu}\phi$
$-\frac{1}{4}\partial_{\omega}\phi\partial^{\omega}\phi\widetilde{g}_{\mu\nu}-V(\phi)\widetilde{g}_{\mu\nu}\Big{]}\delta\widetilde{g}^{\mu\nu}~{}~{}.$
Variation of the fourth term of the action (27):
$\delta\left(\frac{1}{2}\partial_{[e}Z_{ad]}\partial^{[e}Z^{ad]}\widetilde{g}^{ac}\widetilde{g}^{be}\widetilde{g}^{fd}\sqrt{-\widetilde{g}}\right)$
$=\frac{3}{2}\sqrt{-\widetilde{g}}\partial_{[e]}Z_{a\mu]}\partial^{[e]}Z^{a\sigma]}\widetilde{g}_{\sigma\nu}\delta\widetilde{g}^{\mu\nu}$
$-\frac{1}{2}\delta\widetilde{g}^{\mu\nu}\sqrt{-\widetilde{g}}\partial_{[e}Z_{ad]}\partial^{[e}Z^{ad]}~{}~{}.$
Variation of the fifth (or sixth) term of the action (27):
$\delta\left(-\kappa\sqrt{-\widetilde{g}}\widetilde{g}^{ac}\widetilde{g}^{be}\widetilde{g}^{fd}\partial_{[e}\phi Z_{ad]}\partial^{[e}Z^{ad]}\right)$
$=\frac{\kappa}{2}\sqrt{-\widetilde{g}}\widetilde{g}_{\mu\nu}\delta\widetilde{g}^{\mu\nu}\partial_{[e}\phi Z_{ad]}\partial^{[e}Z^{ad]}$
$-3\kappa\sqrt{-\widetilde{g}}\widetilde{g}_{\sigma\nu}\delta\widetilde{g}^{\mu\nu}\partial_{[e}\phi Z_{a\mu]}\partial^{[e}Z^{a\sigma]}~{}~{}.$
Variation of the seventh term of the action (27):
$\delta\left(\frac{\kappa^{2}}{2}\sqrt{-\widetilde{g}}\widetilde{g}^{ac}\widetilde{g}^{be}\widetilde{g}^{fd}\partial_{[e}\phi~{}Z_{ad]}\partial_{[f}\phi~{}Z_{bc]}\right)$
$=3\frac{\kappa^{2}}{2}\sqrt{-\widetilde{g}}\partial_{[e}\phi~{}Z_{a\mu]}\partial^{[e}\phi~{}Z^{a\sigma]}\widetilde{g}_{\sigma\nu}\delta\widetilde{g}^{\mu\nu}$
$-\frac{\kappa^{2}}{4}\sqrt{-\widetilde{g}}\partial_{[e}\phi~{}Z_{ad]}\partial^{[e}\phi~{}Z^{ad]}\widetilde{g}_{\mu\nu}\delta\widetilde{g}^{\mu\nu}~{}~{}.$
Using all these terms now we can apply the least action principle to obtain the gravitational equation of motion corresponding to the Einstein action (27), and is given by:
$\begin{split}\frac{1}{2\kappa^{2}}\widetilde{G}_{\mu\nu}&=\frac{1}{2}\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{4}\partial_{\omega}\phi\partial^{\omega}\phi\widetilde{g}_{\mu\nu}-\frac{3}{2}\partial_{[e}Z_{a\mu]}\partial^{[e}Z^{a\sigma]}\widetilde{g}_{\sigma\nu}\\ &+\frac{1}{2}\widetilde{g}_{\mu\nu}\partial_{[e}Z_{ad]}\partial^{[e}Z^{ad]}-\frac{\kappa}{2}\widetilde{g}_{\mu\nu}\partial_{[e}\phi Z_{ad]}\partial^{[e}Z^{ad]}\\ &+3\kappa\widetilde{g}_{\sigma\nu}\partial_{[e}\phi Z_{a\mu]}\partial^{[e}Z^{a\sigma]}+\frac{\kappa^{2}}{4}\partial_{[e}\phi~{}Z_{ad]}\partial^{[e}\phi~{}Z^{ad]}\widetilde{g}_{\mu\nu}\\ &-3\frac{\kappa^{2}}{2}\partial_{[e}\phi~{}Z_{a\mu]}\partial^{[e}\phi~{}Z^{a\sigma]}\widetilde{g}_{\sigma\nu}+\frac{V(\phi)\widetilde{g}_{\mu\nu}}{2}~{}~{}.\end{split}$ (31)
Eq. (31) results to the Friedmann Eq. (29) and Eq. (30) with the consideration of a spatially flat FLRW metric.